Solve the following equation
2x – 5 (4-x) = 3(2x – 6) + x – 2

Answer by checkley71(8403)

2X-5(4-X)=3(2X-6)+X-2 OR 2X-20+5X=6X-18+X-2 OR 0=0 OR 7X-20=7X-20 OR
7X-7X=20-20 OR 0=0